# Generalized Multi-Objective Reinforcement Learning with Envelope Updates in URLLC-enabled Vehicular Networks

### v2

@@ Abstract
We develop a novel multi-objective reinforcement learning (MORL) framework to jointly optimize wireless network selection and autonomous driving policies in a multi-band vehicular network operating on conventional sub-6GHz spectrum and Terahertz frequencies. The proposed framework is designed to (i) maximize the traffic flow and minimize collisions by controlling the vehicle’s motion dynamics (i.e., speed and acceleration), and (ii) enhance the ultra-reliable low-latency communication (URLLC) while minimizing handoffs (HOs). We cast this problem as a multi-objective Markov Decision Process (MOMDP) and develop solutions for both predefined and unknown preferences of the conflicting objectives. Specifically, deep-Q-network and double deep-Q-network-based solutions are developed first that consider scalarizing the transportation and telecommunication rewards using predefined preferences. We then develop a novel envelope MORL solution which develop policies that address multiple objectives with unknown preferences to the agent. While this approach reduces reliance on scalar rewards, policy effectiveness varying with different preferences is a challenge. To address this, we apply a generalized version of the Bellman equation and optimize the convex envelope of multi-objective Q values to learn a unified parametric representation capable of generating optimal policies across all possible preference configurations. Following an initial learning phase, our agent can execute optimal policies under any specified preference or infer preferences from minimal data samples. Numerical results validate the efficacy of the envelope-based MORL solution and demonstrate interesting insights related to the inter-dependency of vehicle motion dynamics, HOs, and the communication data rate. The proposed policies enable autonomous vehicles (AVs) to adopt safe driving behaviors with improved connectivity.
@@ I Introduction
On the other hand, the use of Deep Reinforcement Learning (DRL) is becoming critical for online decision making in highly random, mobile-oriented wireless environments. In the context of V2I communications, a wealth of research has focused on improving network quality of service (QoS) (e.g., including transmission delay, link throughput, etc.) via DRL-based resource allocation [2, 3, 4]. This research has focused on considering subchannel and power allocation to improve V2I communication. In particular, the authors in [2, 3] adopted deep-Q network (DQN) and multi-agent DQN to simultaneously improve the overall throughput of V2I links and the payload delivery rate of vehicle-to-vehicle (V2V) links. Xu et al. [4] derived the contribution-based dual-clip proximal policy to optimise V2I and V2V connections separately. However, their system model includes only a single BS where handovers (HOs) are not considered.
@@
Along another note, most of the existing research works in the transportation are focused on collision-avoidance [10, 11], safe driving [12], and efficient fuel consumption [12, 13, 14]. For instance, in [10], the authors applied a RL framework for faster travel and the reward is proportional to the AV’s velocity along with a penalty for vehicle collision. The action space included acceleration, deceleration, lane changes (LC), and maintain speed, whereas the state space was based on AVs’ locations and their respective velocities. In [12], the authors applied DDQN to enhance AVs’ driving safety and fuel consumption. The state space included AVs’ locations, fuel consumption, and velocities, whereas the actions included speeds and LC of AVs. In [13], the authors applied the Intelligent Driver Model (IDM) and Minimizing Overall Braking Induced by Lane Change (MOBIL) to control steering and lane change. The proposed reward design encourages long traffic, high speed and discourages unnecessary LC. Authors in [14] introduced multi-objective actor-critic to improve the tradeoff between energy consumption and travel efficiency of AVs. The authors derived MO actor-critic to optimize two objectives. Recently, a few optimization-based research works have started looking into this problem where the AV velocity has been optimized using simplified car-following models [15, 16, 17, 18, 19, 20], but these models are computationally expensive, do not learn from past experiences thus not relevant for faster decision-making, and are not applicable to dynamic traffic scenarios with various control and communication parameters.
@@
The rest of this work is organized as follows. Section II shows the system model, and Section III illustrates the RL, DRL, MORL background, and problem formulation. Section IV introduces the Multi-Objective Reinforcement Learning Approach. The simulations are presented in Section V, and Section VI concludes this research work.
@@ II-A Downlink V2I Data Transmission Model
The signal transmitted by the RBS is subject to path-loss and short-term channel fading. Subsequently, the signal-to-interference-plus-noise ratio (SINR) of the $j$-th AV from $i$-th RBS is given as [25]:
@@
In the context of a Terahertz (THz) network, where molecular absorption significantly impacts signal propagation, the significance of line-of-sight (LOS) transmissions over non-line-of-sight (NLOS) transmissions is dominant. Consequently, the SINR for a given $j$-th AV can be modeled as follows [25]:
@@ II-B Transportation Model
III Preliminaries and MOMDP Formulation
This section first provides a primer on the formulation of the Markov Decision Process (MDP) for single-objective problems and the Multi-Objective Markov Decision Process (MOMDPs) for multi-objective problems. We then formulate our considered problem as an MOMDP and discuss the design of state-action space and rewards function.
III-A Mathematical Preliminaries
III-A1 MDPs
In the distribution of tasks pool, the interaction between tasks and agents can be defined as MDP $\mathcal{M}$, represented by a tuple, $\mathcal{M}=<\mathcal{S},\mathcal{A},r,\mathcal{P},\mathcal{\gamma}>$, where $\mathcal{S},\mathcal{A},r,$ are the set of states $s$, actions $a$, and reward $r(s,a)$, respectively. $\gamma\in[0,1)$ is the discount factor. $r(s,a)$ represents the stochastic instantaneous reward value that the agent can receive given a specific action $a\in\mathcal{A}$ has been taken in a given specific state $s\in\mathcal{S}$. $\mathcal{P}(s_{t+1}|s_{t},a_{t})$ indicates the transition probability for the agent to take an action $a_{t}\in\mathcal{A}$ on state $s_{t}\in\mathcal{S}$ to the next state $s_{t+1}\in\mathcal{S}$ in time step $t$. In RL, the agent interacts with the environment by following a trajectory $\tau=\{(s_{t},a_{t})\}_{t=0}^{\infty}$ from time stamp $t=0$ to the end of the interacting episode, and receives a total discounted reward as ${r}_{\tau}=\sum_{t=0}^{\infty}\mathcal{\gamma}\cdot r_{t}(s_{t},a_{t})$. The goal in RL is to find the mapping policy $\pi$ between states and actions that maximizes the discounted total reward. In each time step $t$, the agent selects an action based on its current state to maximize the discounted total reward ${r}_{\tau}$. Finally, we can obtain a policy $\pi$ that belongs to the policy set $\Pi$. The state-action value function $Q_{\pi}(s,a)$ of the policy $\pi$ on state $s$ can be given as:
$Q_{\pi}(s,a)=\mathbb{E}_{\pi}\left[r_{t}(s_{t},a_{t})+\mathcal{\gamma}Q_{\pi}(s_{t-1},a_{t-1})\right]$ (13)
Thus, the optimal policy is given as $\pi^{*}=\max_{\pi\in\Pi}Q_{\pi}(s,a).$
III-A2 MOMDPs
The goal of MORL is to obtain policies among $M$ conflicting or competing objectives, where the relative importance (preferences) of each objective may be known or unknown to the agent. Similar to RL, MORL can be formulated by MOMDP which extends the MDP by defining a new reward space, preference space, and preference function, i.e., $\mathcal{M}=<\mathcal{S},\mathcal{A},\mathbf{r},\mathcal{P},\Omega,\iota_{0}>$, where $\mathbf{r}\in\mathbb{R}^{H}$ is a vector of reward functions corresponding to $J$ objectives, e.g., $\mathbf{r}=[r^{1},r^{2},\dots,r^{H}]$, $\Omega$ is the preference space where $\boldsymbol{\omega}\in\Omega$ is the preference vector corresponding to $H$ objectives, and $\sum_{h=1}^{H}{{\omega}}^{h}=1$. $\iota_{0}$ is the probability distribution over initial states. In MOMDPs, a policy $\pi:\mathcal{S}\to\mathcal{A}$ defines a mapping from states to actions with the goal of maximizing a vector of expected rewards. Given the distribution $\iota_{0}$ and a policy $\pi$, the expected discounted return is defined as:
@@ III-A2 MOMDPs
Considering all returns from MOMDP, we have Pareto frontier set :=$\mathcal{F}^{*}:=\left\{\mathbf{\hat{r}}\mid\nexists\mathbf{\hat{r}}^{\prime}\geq\mathbf{\hat{r}}\right\}$ [31], where $\mathbf{\hat{r}}=\sum_{t=0}^{\infty}\gamma\cdot\mathbf{r}(s_{t},a_{t})$. For all possible preferences in $\Omega$, we define a convex coverage set (CCS) of the Pareto frontier which contains all returns that provide the maximum cumulative reward, i.e.,
@@
III-B MOMDP Formulation
In this section, we formulate the considered problem as a MOMDP and discuss the corresponding design of state, action, and reward. The state transitions and rewards are a function of the AV environment and actions taken by the AV.
III-B1 State Space
@@
III-B2 Two Dimensional Action Space
@@ III-B3 Rewards
IV Proposed Single-Policy and Multi-Policy MORL Algorithms
In contrast to conventional DRL, MORL requires the agent to optimize multiple objectives simultaneously. These objectives might have predefined preferences, or the preferences could be unknown. In this section, we first investigate the single policy solutions to the MORL problem with predefined preferences, as discussed in Section IV-A. Subsequently, a multiple-policy envelope solution for MORL is proposed in Section IV for cases where the preferences are unknown.
IV-A MORL Solution with Predefined Preferences
Given a set of preferences in MORL problems, single policy algorithms aim to scalarize the reward value to determine the best policy, considering the relative priorities assigned to competing objectives. We explore two DRL methods: DQN [1] and DDQN [32] for MORL, each of which employs a neural network parameterized by $\boldsymbol{\theta}_{t}$ (in each time step $t$) to approximate the $Q$-value function for a state-action pair, i.e., $Q(s_{t},a_{t};\boldsymbol{\theta}_{t})$. After taking action $a_{t}$ in state $s_{t}$ and receiving instant reward $r_{t+1}$, we can formulate a target $Q$ function ${\hat{Q}}(s_{t},a_{t})$ as in Eq. (13), which is used to optimize the neural network $\boldsymbol{\theta}_{t}$ using gradient descent, as given in [32],
IV-A1 Multi-Objective Deep Q-Network (MO-DQN)
The proposed MO-DQN method incorporates a target Q-network and experience replay to stabilize the learning process and ensure convergence, as discussed in the following:
  • Experience Replay: To address issues related to correlations between sequential observations and to improve data efficiency, MO-DQN utilizes the experience replay mechanism, which stores past transition tuples $(s_{z},a_{z},s_{z+1},r_{z})$ in a replay buffer $\mathcal{D}_{\mathcal{T}}$ with size $N_{\mathcal{T}}$, i.e., $z\in\{1,\dots,N_{\mathcal{T}}\}$. During the training phase, mini-batches of these transitions are randomly sampled from the buffer. This method not only reduces the variance of each update but also allows the neural network to benefit from learning across a diverse range of past experiences, thus avoiding local optima and overfitting [34, 35].
IV-A2 Multi-Objective Double DQN (MO-DDQN)
Although single-policy methods are adequate when we possess prior knowledge of task preferences, the acquired policy is constrained in its adaptability to situations with varying preferences. For instance, collision avoidance may not remain a high priority in highway environments with reduced vehicle density. Also, in traffic jams or parking lots where AVs are still, the preference for telecommunication rewards becomes higher. In the next section, we seek to design the multi-policy algorithm that handles unknown preferences in multi-objective RL scenarios.
Initialization:
Initialize $Q$-network weights $\boldsymbol{\theta}$ randomly,
while episode $<$ episode limit and runtime $<$ time limit do
       Initialize $t\leftarrow 0$ and state $s_{t}$ based on environment
       while $t\leq T_{hl}$  do
             Derive $a^{\mathrm{tran}}_{t}$ and $a^{\mathrm{tele}}_{t}$ from $a_{t}$
             Apply $a^{\text{tran}}_{t}$ and $a^{\text{tele}}_{t}$,
             observe reward $r_{t}$ and next state $s_{t+1}$.
             Store transition $(s_{t},a_{t},s_{t+1},r_{t})$ in $\mathcal{D}_{\mathcal{T}}$.
             where $z\in\{1,\ldots,N_{\mathcal{T}}\}$.
             Set target-$Q$ for each sampled transition:
             for each transition $z$ do
                   if episode ends at step $z+1$ then
                         $\hat{Q}(s_{z},a_{z})=r_{z}$
                   else
                   end if
             end for
             if MO-DDQN then
             end if
            $t\leftarrow t+1$
       end while
      Update policy $\pi$ based on learned $Q$.
end while
Algorithm 1 Multi-Objective Double Deep Q-Learning
IV-B Multi-Policy Envelope MORL Algorithm
@@
Our proposed MO-DDQN-Envelope algorithm is designed to learn a spectrum of Pareto-optimal policies simultaneously within a preference space ${\Omega}$, as described in Section IV-A and illustrated in Fig. 3. Unlike the Envelope-MOQ model proposed in [31], which employs the REINFORCE algorithm, our MO-DDQN-Envelope algorithm incorporates DDQN to enhance both convergence stability and sample training efficiency. REINFORCE, as a policy gradient method, often suffers from high variance in complex environments like RF-Thz-Highway-Env, leading to unstable updates. In contrast, DDQN leverages temporal difference (TD) learning, which reduces variance and mitigates the overestimation bias. Furthermore, DDQN employs replay experience to enable the agent to learn from past experiences multiple times, improving sample efficiency, and uses target networks to stabilize learning by ensuring consistent targets during updates.
@@
IV-B1 Bellman Operation with Optimal Filter
@@
Unlike scalarized single-policy approaches discussed in Section IV-A, which fail to adapt the scalar utility across varying $\boldsymbol{\omega}$, the convex envelope formulation explicitly leverages the supremum operator to optimize across all possible actions $a$ and preference vectors $\boldsymbol{\omega}^{\prime}$.
IV-B2 Hindsight Experience Replay (HER)
@@ IV-B3 Homotopy Optimization
$\end{multlined}\hat{\mathbf{Q}}(s_{t},a_{t},\boldsymbol{\omega}_{t};\boldsymbol{\theta}_{t}^{\prime})=\mathbb{E}_{s_{t+1}}[\mathbf{r}_{t}+\gamma\arg_{Q}\max_{a_{t},\boldsymbol{\omega}_{t}^{\prime}}\boldsymbol{\omega}^{T}\mathbf{Q}(s_{t+1},a_{t},\boldsymbol{\omega}_{t}^{\prime};\boldsymbol{\theta}_{t}^{\prime})]$ (26)
@@
IV-C Complexity Analysis
IV-C1 Training Complexity Analysis
@@
IV-C2 Evaluation Complexity Analysis
@@ V-A Simulation Environment
We utilize a Multi-Layer Perceptron (MLP) to construct the $Q_{\theta}$ neural network for the training and evaluation phases. The architecture consists of three fully connected layers (FNNs), each containing 128 neurons. Following these layers, we apply the ReLU activation function to improve training efficiency. The output layer of the target policy network $Q_{\theta}^{\prime}$ employs a sigmoid activation function to constrain the output range of actions. This ensures that the action values remain within a feasible and interpretable range, which is critical for MOO decision-making tasks. This architecture was chosen after a comprehensive hyperparameter tuning process. Simulation parameters are detailed in Table II unless otherwise specified.
@@
We evaluated all methods across five test instances, varying key variables: desired minimum and maximum longitudinal velocities ($v_{\text{min}},v_{\text{max}}$), the number of TBSs ($n_{T}$), and the number of AVs ($M$). These combinations are detailed in Table III.

### v3

@@ Abstract
We develop a novel multi-objective reinforcement learning (MORL) framework to jointly optimize wireless network selection and autonomous driving policies in a multi-band vehicular network operating on conventional sub-6GHz spectrum and Terahertz frequencies. The proposed framework is designed to (i) maximize the traffic flow and minimize collisions by controlling the vehicle’s motion dynamics (i.e., speed and acceleration), and (ii) enhance the ultra-reliable low-latency communication (URLLC) while minimizing handoffs (HOs). We cast this problem as a multi-objective Markov Decision Process (MOMDP) and develop solutions for both predefined and unknown preferences of the conflicting objectives. Specifically, we develop a novel envelope MORL solution which develops policies that address multiple objectives with unknown preferences to the agent. While this approach reduces reliance on scalar rewards, policy effectiveness varying with different preferences is a challenge. To address this, we apply a generalized version of the Bellman equation and optimize the convex envelope of multi-objective Q values to learn a unified parametric representation capable of generating optimal policies across all possible preference configurations. Following an initial learning phase, our agent can execute optimal policies under any specified preference or infer preferences from minimal data samples. Numerical results validate the efficacy of the envelope-based MORL solution and demonstrate interesting insights related to the inter-dependency of vehicle motion dynamics, HOs, and the communication data rate. The proposed policies enable autonomous vehicles (AVs) to adopt safe driving behaviors with improved connectivity.
@@ I Introduction
On the other hand, the use of Deep Reinforcement Learning (DRL) is becoming critical for online decision making in highly random, mobile-oriented wireless environments. In the context of V2I communications, a wealth of research has focused on improving network quality of service (QoS) (e.g., including transmission delay, link throughput, etc.) via DRL-based resource allocation [2, 3, 4]. This research has focused on considering subchannel and power allocation to improve V2I communication. In particular, the authors in [2, 3, 5] adopted deep-Q network (DQN) and multi-agent DQN to simultaneously improve the overall throughput of V2I links and the payload delivery rate of vehicle-to-vehicle (V2V) links. Xu et al. [4] derived the contribution-based dual-clip proximal policy to optimise V2I and V2V connections separately. However, their system model includes only a single BS where handovers (HOs) are not considered.
@@
Along another note, most of the existing research works in the transportation are focused on collision-avoidance [11, 12], safe driving [13, 5], and efficient fuel consumption [13, 14, 15]. For instance, in [11], the authors applied a RL framework for faster travel and the reward is proportional to the AV’s velocity along with a penalty for vehicle collision. The action space included acceleration, deceleration, lane changes (LC), and maintain speed, whereas the state space was based on AVs’ locations and their respective velocities. In [13], the authors applied DDQN to enhance AVs’ driving safety and fuel consumption. The state space included AVs’ locations, fuel consumption, and velocities, whereas the actions included speeds and LC of AVs. In [14], the authors applied the Intelligent Driver Model (IDM) and Minimizing Overall Braking Induced by Lane Change (MOBIL) to control steering and lane change. The proposed reward design encourages long traffic, high speed and discourages unnecessary LC. Authors in [15] introduced multi-objective actor-critic to improve the tradeoff between energy consumption and travel efficiency of AVs. The authors derived MO actor-critic to optimize two objectives. Recently, a few optimization-based research works have started looking into this problem where the AV velocity has been optimized using simplified car-following models [16, 17, 18, 19, 20], but these models are computationally expensive, do not learn from past experiences thus not relevant for faster decision-making, and are not applicable to dynamic traffic scenarios with various control and communication parameters.
@@
The rest of this work is organized as follows. Section II shows the system model, and Section III provides MOMDP formulation. Section IV introduces the proposed solution. The simulations are presented in Section V, and Section VI concludes this research work.
@@ II System Model and Assumptions
Compared to the Gipps’ model [25] and learning-based approaches [26], we adopt a transportation model that combines Kinematics + IDM + MOBIL, which is widely used in autonomous driving decision-making, as in [14]. This hybrid model provides improved computational efficiency and adaptability to MORL, outperforming the Gipps and learning-based models in these aspects. For the telecommunication model, in contrast to RF-only 5G V2I, THz-only 6G V2I [1], and RIS-assisted models [27], we adopt a hybrid multiband RF–THz V2I framework [28]. This model offers superior performance in terms of throughput, latency, reliability, mobility management, and HOs control.
@@ II-A Downlink V2I Data Transmission Model
The signal transmitted by the RBS is subject to path-loss and short-term channel fading. Subsequently, the signal-to-interference-plus-noise ratio (SINR) of the $j$-th AV from $i$-th RBS is given as [29, 28]:
@@
In the context of a Terahertz (THz) network, where molecular absorption significantly impacts signal propagation, the significance of line-of-sight (LOS) transmissions over non-line-of-sight (NLOS) transmissions is dominant. Consequently, the SINR for a given $j$-th AV can be modeled as follows:
@@ II-B Transportation Model
III MOMDP Formulation
This section formulates the multi-objective Markov Decision Process (MOMDPs) for the considered problem. We then discuss the design of state-action space and rewards function. The state transitions and rewards are a function of the AV environment and actions taken by the AV.
III-A MOMDPs and Pareto Front
The goal of MORL is to obtain policies among $M$ conflicting or competing objectives, where the relative importance (preferences) of each objective may be known or unknown to the agent. Similar to RL, MORL can be formulated by MOMDP which extends the MDP by defining a new reward space, preference space, and preference function, i.e., $\mathcal{M}=<\mathcal{S},\mathcal{A},\mathbf{r},\mathcal{P},\Omega,\iota_{0}>$, where $\mathbf{r}\in\mathbb{R}^{H}$ is a vector of reward functions corresponding to $H$ objectives, e.g., $\mathbf{r}=[r^{1},r^{2},\dots,r^{H}]$, $\Omega$ is the preference space where $\boldsymbol{\omega}\in\Omega$ is the preference vector corresponding to $H$ objectives, and $\sum_{h=1}^{H}{{\omega}}^{h}=1$. $\iota_{0}$ is the probability distribution over initial states. In MOMDPs, a policy $\pi:\mathcal{S}\to\mathcal{A}$ defines a mapping from states to actions with the goal of maximizing a vector of expected rewards. $\mathcal{P}(s_{t+1}|s_{t},a_{t})$ indicates the transition probability for the agent to take an action $a_{t}\in\mathcal{A}$ on state $s_{t}\in\mathcal{S}$ to the next state $s_{t+1}\in\mathcal{S}$ in time step $t$. Given the distribution $\iota_{0}$ and a policy $\pi$, the expected discounted return is defined as:
@@ III-A MOMDPs and Pareto Front
Considering all returns from MOMDP, we have Pareto frontier set :=$\mathcal{F}^{*}:=\left\{\mathbf{\hat{r}}\mid\nexists\>\mathbf{\hat{r}}^{\prime}\geq\mathbf{\hat{r}}\right\}$ [35], where $\mathbf{\hat{r}}=\sum_{t=0}^{\infty}\gamma\cdot\mathbf{r}(s_{t},a_{t})$. For all possible preferences in $\Omega$, we define a convex coverage set (CCS) of the Pareto frontier which contains all returns that provide the maximum cumulative reward, i.e.,
@@
III-B State Space
@@
III-C Two Dimensional Action Space
@@ III-D Rewards
IV Multi-Policy Envelope MORL Algorithm
In contrast to conventional DRL, MORL requires the agent to optimize multiple objectives simultaneously. These objectives might have predefined preferences, or the preferences could be unknown. A multiple-policy envelope solution for MORL is proposed in Section IV for unknown preferences. The single policy solutions to the MORL problem with predefined preferences are discussed in Appendix. Although single-policy methods are adequate when we possess prior knowledge of task preferences, the acquired policy is constrained in its adaptability to situations with varying preferences. For instance, collision avoidance may not remain a high priority in highway environments with reduced vehicle density. Also, in traffic jams or parking lots where AVs are still, the preference for telecommunication rewards becomes higher.
Our approach separates the model training and evaluation phases. The MORL agent is trained offline using a predefined simulated environment. During the evaluation stage, we use a pool of pretrained MORL models, each tailored to a predefined configuration defined by a given number of AVs (representing traffic conditions), number of RF/THz BSs, and a specific topology. This approach is also adopted by 3GPP [36], i.e., at run time, the controller simply selects the MORL model that best matches the observed snapshot, achieving zero online training overhead [36].
@@ IV Multi-Policy Envelope MORL Algorithm
Our proposed MO-DDQN-Envelope algorithm is designed to learn a spectrum of Pareto-optimal policies simultaneously within a preference space ${\Omega}$, as described in Section III-A and illustrated in Fig. 3. Unlike the Envelope-MOQ model proposed in [35], which employs the REINFORCE algorithm, our MO-DDQN-Envelope algorithm incorporates DDQN to enhance both convergence stability and sample training efficiency. REINFORCE, as a policy gradient method, often suffers from high variance in complex environments like RF-Thz-Highway-Env, leading to unstable updates. In contrast, DDQN leverages temporal difference (TD) learning, which reduces variance and mitigates the overestimation bias. Furthermore, DDQN employs replay experience to enable the agent to learn from past experiences multiple times, improving sample efficiency, and uses target networks to stabilize learning by ensuring consistent targets during updates.
@@
IV-1 Bellman Operation with Optimal Filter
@@
Unlike scalarized single-policy approaches discussed in Appendix, which fail to adapt the scalar utility across varying $\boldsymbol{\omega}$, the convex envelope formulation explicitly leverages the supremum operator to optimize across all possible actions $a$ and preference vectors $\boldsymbol{\omega}^{\prime}$.
IV-2 Hindsight Experience Replay (HER)
@@ IV-3 Homotopy Optimization
[…]
@@
IV-A Complexity Analysis
IV-A1 Training Complexity Analysis
@@
IV-A2 Evaluation Complexity Analysis
@@ V-A Simulation Environment
We utilize a Multi-Layer Perceptron (MLP) to construct the $Q_{\theta}$ neural network for the training and evaluation phases. The architecture consists of three fully connected layers (FNNs), each containing 128 neurons. Following these layers, we apply the ReLU activation function to improve training efficiency. The output layer of the target policy network $Q_{\theta}^{\prime}$ employs a sigmoid activation function to constrain the output range of actions. This ensures that the action values remain within a feasible and interpretable range, which is critical for MOO decision-making tasks. This architecture was chosen after a comprehensive hyperparameter tuning process. This configuration strikes a balance between model accuracy, training efficiency, and convergence performance. Simulation parameters are detailed in Table II unless otherwise specified.
@@
We evaluated all methods across five test instances, varying key variables: desired minimum and maximum longitudinal velocities ($v_{\text{min}},v_{\text{max}}$), the number of TBSs ($n_{T}$), and the number of AVs ($M$). These combinations are given in Table III.